Find the height or length of these natural wonders in kilometers, meters, and centimeters (a) a cave system with a mapped length of 326 miles km cm (b) a waterfall that drops 1,043.2 ft km cm (c) a 17,000 ft tall mountain km cm (d) a canyon with a depth of 65,200 ft km cm Need Help? Readt
Answer #1

(a) 1 miles = 1.609 km

326 miles = 326*1.609 = 524.534 km​​​​​​

1 km = 1000 m

326 miles = 524.534 km = 5.245*10^5 m

1 m = 100 cm

326 miles = 524.534 km = 5.245*10^5 m = 5.245*10^7 cm

(b) 1 feet = 0.000305 km

1043.2 ft = 1043.2 * 0.000305 km = 0.318176 km​​​​​​

1043.2 ft = 0.318176 km = 318.176 m = 3.18*10^4 cm